A sample of compressed oxygen has a volume of 50L and a pressure of 140atm at 25oC. The cylinder is cooled until the pressure is 10atm. What is the new temperature in the cylinder?


1
Expert's answer
2020-09-29T06:12:23-0400

Question # 133172

A sample of compressed oxygen has a volume of 50L and a pressure of 140atm at 25oC. The cylinder is cooled until the pressure is 10atm. What is the new temperature in the cylinder?

Answer

P1V1/T1 =P2V2/T2

P1= initial pressure, V1= initial volume, T1= initial temperature in kelvins

P2= final pressure, V2= final volume, T2= final temperature in kelvins

P1V1/T1 =P2V2/T2

(140atm×50L)/323K= (10atm×50L)/ T2

T2 = (323K×10atm×50L) ÷ (140atm×50L)

               = 23.07 K